### v2

# A Bias-Correction Decentralized Stochastic Gradient Algorithm with Momentum Acceleration

Abstract
Distributed stochastic optimization algorithms can process large-scale datasets simultaneously, significantly accelerating model training. However, their effectiveness is often hindered by the sparsity of distributed networks and data heterogeneity. In this paper, we propose a momentum-accelerated distributed stochastic gradient algorithm, termed Exact-Diffusion with Momentum (EDM), which mitigates the bias from data heterogeneity and incorporates momentum techniques commonly used in deep learning to enhance convergence rate. Our theoretical analysis demonstrates that the EDM algorithm converges sub-linearly to the neighborhood of the optimal solution, the radius of which is irrespective of data heterogeneity, when applied to non-convex objective functions; under the Polyak-Łojasiewicz condition, which is a weaker assumption than strong convexity, it converges linearly to the target region. Our analysis techniques employed to handle momentum in complex distributed parameter update structures yield a sufficiently tight convergence upper bound, offering a new perspective for the theoretical analysis of other momentum-based distributed algorithms.
1 Introduction
The distributed stochastic optimization (DSO) problem has garnered significant attention across various domains, including the Internet of Things (IoT) (Ali et al., 2018), federated learning (Beltrán et al., 2023), and statistical machine learning (Boyd et al., 2011). In this paper, we specifically address the following optimization problem:
$\mathop{\min}\limits_{\mathbf{x}\in\mathbb{R}^{d}}\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x}),\quad f_{i}(\mathbf{x})\triangleq\mathbb{E}_{\bm{\xi}_{i}\sim\mathcal{D}_{i}}F_{i}(\mathbf{x},\bm{\xi}_{i}),$ (1.1)
where $F_{i}(\mathbf{x},\bm{\xi}_{i})$ represents the loss function at agent $i$ given the parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $\bm{\xi}_{i}$ denotes the data on agent $i$, which is independently and identically generated from the distribution $\mathcal{D}_{i}$. Each agent is capable of exchanging information with its neighbors, thus facilitating collaborative decision-making to optimize the overall global objective function.
A common approach to solving DSO problems is through first-order methods (Xin et al., 2020). Notably, Nedic and Ozdaglar (2009) proposed a sub-gradient optimization algorithm for distributed optimization (DO), which serves as the deterministic counterpart of DSO. By leveraging the concept of stochastic approximation, these algorithms can be effectively adapted to tackle distributed stochastic optimization problems by estimating the gradient corresponding to the current parameters of each agent. Jiang et al. (2017) and Lian et al. (2017) introduced a stochastic variant of Yuan et al. (2016), referred to as Distributed Stochastic Gradient Descent (DSGD), and investigated its convergence properties for non-convex objective functions. Following this, a series of improved stochastic optimization algorithms were developed (Spiridonoff et al., 2020; Zhang and You, 2019; Tang et al., 2018).
In a distributed scenario, data heterogeneity among agents presents a significant challenge (Hsieh et al., 2020). Several works Yuan et al. (2021); Lin et al. (2021) have indicated that DSGD can converge to a neighborhood of the optimal value, with the radius of this neighborhood determined by the variance of the stochastic gradient and the heterogeneity of the data for both strongly convex optimization and non-convex optimization. The sparsity of the network further exacerbates the impact of heterogeneity (Koloskova et al., 2020; Yuan et al., 2020). A viable strategy to mitigate this issue is to modify the gradient update to approximate the average gradient, collectively referred to as bias-correction algorithms. For instance, Zhang and You (2019) proposed Distributed Stochastic Gradient Tracking (DSGT), a widely used bias-correction algorithm, and demonstrated its sub-linear convergence for non-convex objective functions. Its convergence properties in the strongly convex case were established in Pu and Nedić (2021). In contrast, exact-diffusion (Yuan et al., 2020), also denoted as D2 in Tang et al. (2018), effectively eliminates the effects of data heterogeneity without requiring additional information aggregation. A unified analysis of the bias-correction algorithms is conducted in Alghunaim and Yuan (2022), and Alghunaim and Yuan (2024) noted that the convergence rate of DSGT is slightly inferior to ED/D2.
To enhance convergence properties, a common strategy is to incorporate the momentum method (Nesterov, 2013; Polyak, 1964). Works by Yu et al. (2019) and Gao and Huang (2020) have introduced momentum into distributed stochastic optimization problems, providing convergence rates in the non-convex case. However, the introduction of momentum does not eliminate heterogeneous bias, which can render the convergence results less robust Yuan et al. (2021); Takezawa et al. (2023). Some methods, such as DecentLaM Yuan et al. (2021) and Quasi-Global Lin et al. (2021), partially address this issue by modifying the momentum architecture. Introducing momentum into the bias-correction algorithm presents a method that can fundamentally eliminate heterogeneity while retaining the benefits of momentum acceleration. Several studies have explored the incorporation of momentum methods into DSGT and their convergence properties (Takezawa et al., 2023; Gao et al., 2023). Huang et al. (2024) introduced a loopless Chebyshev acceleration (LCA) technique to improve the consensus rate and prove its convergence properties under the PL condition case and general non-convex case. Nonetheless, the momentum version of ED/D2, which functions as a bias-correction algorithm, has received limited attention, despite the theoretical results of Alghunaim and Yuan (2022) indicating that ED/D2 should exhibit superior convergence properties in sparse networks and greater robustness in the presence of highly heterogeneous data.
Regarding the theoretical properties of the momentum-based algorithms, Liu et al. (2020) demonstrated that centralized stochastic gradient descent with momentum possesses the same convergence bounds as standard stochastic gradient descent. However, existing theoretical analyses of momentum-based bias-correction algorithms often overlook this property and impose additional constraints on step size or momentum selection. For instance, Takezawa et al. (2023); Gao et al. (2023) require step size $\alpha=\mathcal{O}((1-\lambda)^{2})$ and Huang et al. (2024) require the momentum parameter $1-\beta=\mathcal{O}(1-\lambda)$. In contrast, Alghunaim and Yuan (2022) only requires a step size of $\alpha=\mathcal{O}(1-\lambda)$. Therefore, whether momentum-based algorithms can achieve convergence rates comparable to those of the original algorithms remains an important question.
Table 1: General Non-Convex Convergence Rate. We focus on the condition when $1-\lambda$ is very small and treat the momentum parameter $\beta$ as a constant. Our conclusion regarding EDM eliminates the impact of the network sparsity $\underline{\lambda}$, the minimum non-zero eigenvalue of $\mathbf{W}$, as stated in Alghunaim and Yuan (2022). Additionally, when $\beta=0$, our conclusion aligns with Yuan et al. (2020). Our conclusion demonstrates that the ED/D2 can eliminate the influence of heterogeneity more rapidly than the DSGT, since the heterogeneous term (the bold part) of EDM is scales with $(1-\lambda)^{-2}$, whereas DSGT scales with $(1-\lambda)^{-3}$.
Method Work Momentum Bound of $\alpha$ Convergence Rate
ED/D2 Alghunaim and Yuan (2022) - $\mathcal{O}(1-\lambda)$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{(1-\lambda)\underline{\lambda}}+\frac{\alpha^{4}L^{4}\sigma^{2}}{(1-\lambda)^{3}\underline{\lambda}}+\bm{\frac{\alpha^{2}L^{2}\zeta_{0}^{2}}{(1-\lambda)^{2}\underline{\lambda}T}}\right)$
Ours Ours $\mathcal{O}(1-\lambda)$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{1-\lambda}+\bm{\frac{\alpha^{2}L^{2}\zeta_{0}^{2}}{(1-\lambda)^{2}T}}\right)$
DSGT Alghunaim and Yuan (2022) - $\mathcal{O}(1-\lambda)$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{1-\lambda}+\frac{\alpha^{4}L^{4}\sigma^{2}}{(1-\lambda)^{4}}+\bm{\frac{\alpha^{2}L^{2}\zeta_{0}^{2}}{(1-\lambda)^{3}T}}\right)$
DSGT-HB Gao et al. (2023) $\mathcal{O}((1-\lambda)^{2})$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{(1-\lambda)^{3}}+\bm{\frac{\alpha^{2}L^{2}\zeta_{0}^{2}}{(1-\lambda)^{3}T}}\right)$
DmSGD Yu et al. (2019) $\mathcal{O}((1-\lambda)^{2})$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{1-\lambda}+\bm{\frac{\alpha^{2}L^{2}\zeta^{2}}{(1-\lambda)^{2}}}\right)$
DecentLaM Yuan et al. (2021) $\mathcal{O}((1-\lambda)^{2})$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{1-\lambda}+\bm{\frac{\alpha^{2}L^{2}\zeta^{2}}{(1-\lambda)^{2}}}\right)$
Quasi-Global Lin et al. (2021) $\mathcal{O}(1-\lambda)$ $\mathcal{O}\left(\frac{1}{\alpha T}+\frac{\alpha L\sigma^{2}}{n}+\frac{\alpha^{2}L^{2}\sigma^{2}}{(1-\lambda)^{2}}+\bm{\frac{\alpha^{2}L^{2}\zeta^{2}}{(1-\lambda)^{2}}}\right)$
The contributions of this paper can be summarized as follows.
  1. 1.
    We propose a novel decentralized optimization algorithm that incorporates momentum into the ED/D2 framework. This algorithm can accelerate the ED/D2 algorithm in decentralized scenarios. And also ensures that our approach will not be affected by data heterogeneity and the network sparsity and will eventually converge to some neighborhood of the optimal solution, thus being more robust to the data heterogeneity. To the best of our knowledge, this work is the first work to investigate the momentum version of ED/D2 and to provide a theoretical analysis. Table 1 shows the theoretical advantages of our algorithm.
  2. 2.
    This paper introduces a new technique to address the variance introduced by the complicated structure of stochastic gradients. Our method establishes a tight upper bound on the convergence rate, which aligns with the findings of the original ED/D2 algorithm as presented in Alghunaim and Yuan (2022), applicable to both non-convex and PL conditions. This alignment reinforces the conclusions drawn by Liu et al. (2020). Furthermore, our proof methodology can be extended to demonstrate the effectiveness of other momentum-based bias-correction algorithms.
This work is arranged as follows. Section 2 addresses the problems to be addressed and the associated challenges. In Section 3, a momentum-based bias-correction algorithm is proposed. Section 4 provides a detailed analysis of the convergence properties of this algorithm. Proofs of relevant theorems are included in the Appendix. The details and results of our simulation are presented in the electronic companion.
2 Preliminaries
2.1 Notations
Throughout this paper, we denote the parameter and data of agent $i$ at time $t$ as $\mathbf{x}_{i}^{(t)}\in\mathbb{R}^{d}$ and $\bm{\xi}_{i}^{(t)}\in\mathbb{R}^{p}$, respectively. Define $\mathbf{X}^{(t)}=(\mathbf{x}_{1}^{(t)},\cdots,\mathbf{x}_{n}^{(t)})^{\top}\in\mathbb{R}^{n\times d}$ and $\bm{\Xi}^{(t)}=(\bm{\xi}_{1}^{(t)},\cdots,\bm{\xi}_{n}^{(t)})^{\top}$. For other variables, we use superscripts to indicate parameters or data at time $t$ and subscripts to denote the labels of agents. We put $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$, $\mathbf{f}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{f}(\mathbf{X})=(f_{1}(\mathbf{x}_{1}),\cdots,f_{n}(\mathbf{x}_{n}))^{\top}\in\mathbb{R}^{n}$. $F_{i}:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}$, $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{F}(\mathbf{X},\bm{\Xi})=(F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$. The gradient of $\mathbf{F}$ w.r.t. $\mathbf{X}$ is defined as $\nabla\mathbf{F}(\mathbf{X},\bm{\Xi})=(\nabla F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,\nabla F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$, and similarly for $\nabla\mathbf{f}(\mathbf{X})$. Let $\mathbf{1}_{d}=(1,\cdots,1)^{\top}\in\mathbb{R}^{d}$ denote the all-ones vector. The mean value of all node parameters is given by $\bar{\mathbf{x}}=\frac{1}{n}\mathbf{X}^{\top}\mathbf{1}_{n}\in\mathbb{R}^{d}$. Additionally, we define the matrix $\overline{\mathbf{X}}=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\mathbf{X}\in\mathbb{R}^{n\times d}$. We define the average of loss functions with the local parameter as $\overline{\mathbf{F}}:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times p}\to\mathbb{R}$, given by $\overline{\mathbf{F}}(\mathbf{X},\bm{\Xi})=\frac{1}{n}\mathbf{F}(\mathbf{X},\bm{\Xi})^{\top}\mathbf{1}_{n}$. The mean of the individual loss functions $f_{i}(\mathbf{x}_{i})$ is denoted as $\bar{\mathbf{f}}:\mathbb{R}^{n\times d}\to\mathbb{R}$ with $\bar{\mathbf{f}}(\mathbf{X})=\frac{1}{n}\mathbf{f}(\mathbf{X})^{\top}\mathbf{1}_{n}$ The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ with $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$. For the momentum, let $\mathbf{m}_{i}^{(t)}$ be the momentum of agent $i$ at time $t$. Let $\mathbf{M}^{(t)}=(\mathbf{m}_{1}^{(t)},\cdots,\mathbf{m}_{n}^{(t)})^{\top}$ and $\bar{\mathbf{m}}^{(t)}=\frac{1}{n}(\mathbf{1}_{n}^{\top}\mathbf{M}^{(t)})^{\top}$. Denote the $\ell_{2}$-norm of $\mathbf{x}\in\mathbb{R}^{d}$ as $\|\mathbf{x}\|$, the Frobenius norm of matrix $\mathbf{A}$ as $\|\mathbf{A}\|_{\mathrm{F}}$ and the spectral norm as $\|\mathbf{A}\|_{\mathrm{op}}$.
2.2 Assumptions
In the decentralized setting, each agent in the network can only communicate with its neighbors to aggregate information. The network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where the set of nodes is $\mathcal{V}=\{1,\cdots,n\}$, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the connections between neighboring nodes. We denote the neighbors of agent $i$ as $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$.
Assumption 1.
Let $\mathbf{W}=(w_{ij})_{1\leq i,j\leq n}$ be the communication matrix corresponding to the network $\mathcal{G}$. The weights $w_{ij}>0$ when $(i,j)\in\mathcal{E}$, and $w_{ij}=0$ otherwise. The matrix $\mathbf{W}$ is designed to satisfy the following conditions:
  1. (1)
    For any $i\in\mathcal{V}$, $w_{ii}>0$;
  2. (2)
    The matrix $\mathbf{W}$ is symmetric and double stochastic, meaning $\mathbf{W}\mathbf{1}_{n}=\mathbf{1}_{n}$ and $\mathbf{1}_{n}^{\top}\mathbf{W}=\mathbf{1}_{n}^{\top}$;
  3. (3)
    The smallest eigenvalue of $\mathbf{W}$ is positive.
Remark 1.
Let $\lambda=\|\mathbf{W}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\|_{\mathrm{op}}$, which is the second largest eigenvalue of $\mathbf{W}$. From (1) and (2) in Assumption 1, we have $\lambda<1$ (Sayed et al., 2014). The quantity $(1-\lambda)^{-1}$ is referred to as the spectral gap of the communication matrix $\mathbf{W}$ (Nedić et al., 2018; Huang et al., 2024), which is related to the efficiency of the consensus among the agents. The relationship between network structure and spectral gap is discussed in detail in Nedić et al. (2018). For example, in a ring graph where each agent is arranged in a circle and can only communicate with its two neighbors, the spectral gap of the network will be $O(n^{2})$. Generally, as the number of agents increases, the network tends to become sparser, which can lead to a larger spectral gap and subsequently hinder the convergence speed.
It is worth noting that (1) and (2) in Assumption 1 also imply that the smallest eigenvalue of $\mathbf{W}$ is greater than $-1$ (Sayed et al., 2014). Therefore, for any symmetric doubly stochastic matrix $\mathbf{W}$, Assumption 1 (3) can be satisfied through the transformation $\widetilde{\mathbf{W}}=\frac{\mathbf{W}+\mathbf{I}}{2}$.
Assumption 2.
(Smooth) Each local objective function $f_{i}(\mathbf{x})$ is $L$-smooth:
$\left\|\nabla f_{i}(\mathbf{x})-\nabla f_{i}(\mathbf{y})\right\|\leq L\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}.$
And each $f_{i}(\mathbf{x})$ is bounded from below, i.e., $f_{i}(\mathbf{x})\geq f_{i}^{\star}=\inf_{\mathbf{x}}f_{i}(\mathbf{x})>-\infty$ for all $\mathbf{x}\in\mathbb{R}^{d}$.
Remark 2.
This assumption is widely utilized in distributed stochastic gradient descent algorithms (Alghunaim and Yuan, 2022; Huang et al., 2024), which mainly restricts the objective function to have a certain smoothness, to ensure that each step of gradient descent is relatively gentle.
Assumption 3.
(Unbiasedness and finite variance) The random data $\bm{\xi}_{i}^{(t)}$ is generated independently and identically from $\mathcal{D}_{i}$. And $\bm{\xi}_{i}^{(t)}$ are independent of each other for all $1\leq i\leq n$ and $t\geq 0$. Let $\mathcal{F}^{(t)}=\mathcal{A}(\bm{\Xi}^{(0)},\cdots,\bm{\Xi}^{(t-1)})$ denote the sigma-algebra generated by $\bm{\Xi}^{(0)},\cdots,\bm{\Xi}^{(t-1)}$. For the $\mathcal{F}^{(t)}$ measurable random vector $\mathbf{x}_{i}^{(t)}$, the following two formula hold,
$\mathbb{E}[\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})-\nabla f_{i}(\mathbf{x}_{i}^{(t)})|\mathcal{F}^{(t)}]=0,$
$\mathbb{E}[\|\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})-f_{i}(\mathbf{x}_{i}^{(t)})\|^{2}|\mathcal{F}^{(t)}]\leq\sigma^{2}.$
Remark 3.
This assumption implies the stochastic gradient at each node is unbiased and possesses a uniform finite variance bound.
Assumption 4.
(Polyak-Łojasiewicz condition) There exists a constant $\mu>0$ such that for all $\mathbf{x}\in\mathbb{R}^{d}$, the objective function $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$ satisfies
$2\mu(f(\mathbf{x})-f^{\star})\leq\left\|\nabla f(\mathbf{x})\right\|^{2}.$
where $f^{\star}$ denotes the optimal value of $f$.
Remark 4.
This assumption is widely used in recent theoretical analyses of decentralized algorithms, e.g. Alghunaim and Yuan (2022); Huang et al. (2024). And the PL-condition is weaker than the $\mu$-strongly convex assumption, which requires $f(\mathbf{y})-f(\mathbf{x})-\langle\nabla f(\mathbf{x}),\mathbf{y}-\mathbf{x}\rangle\geq\frac{\mu}{2}\left\|\mathbf{x}-\mathbf{y}\right\|^{2}$. Many studies on decentralized SGD with momentum acceleration (Lin et al., 2021; Yuan et al., 2021) assume that each agent’s objective function is strongly convex $f_{i}(\mathbf{x})$.
3 Proposed Method
3.1 Drawbacks of momentum-based DSGD algorithm
Decentralized momentum SGD (DmSGD) (Yu et al., 2019; Gao and Huang, 2020) was proposed to accelerate the convergence of distributed learning. In DmSGD, each agent $i$ iteratively updates its momentum and communicates parameters as follows:
$\mathbf{m}_{i}^{(t+1)}$ $=\beta\mathbf{m}_{i}^{(t)}+(1-\beta)\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)}),$ (3.2)
$\mathbf{x}_{i}^{(t+1)}$ $=\sum_{j=1}^{n}w_{ij}\left(\mathbf{x}_{i}^{(t)}-\alpha\mathbf{m}_{i}^{(t+1)}\right).$ (3.3)
However, Proposition 2 in Yuan et al. (2021) demonstrated that DmSGD suffers from an inconsistency bias for strongly convex loss under full-batch gradients (i.e. when $\sigma^{2}=0$),
$\lim_{t\to\infty}\sum_{i=1}^{n}\left\|\mathbf{x}_{i}^{(t)}-\mathbf{x}^{*}\right\|^{2}=\mathcal{O}\left(\frac{\alpha^{2}\zeta^{2}}{(1-\beta)^{2}(1-\lambda)^{2}}\right),$
where $\mathbf{x}^{\star}$ denotes the optimum point, and $\zeta^{2}=(1/n)\sum_{i=1}^{n}\left\|\nabla f_{i}(\mathbf{x}^{*})-\nabla f(\mathbf{x}^{*})\right\|^{2}$ is defined as the measure of the data heterogeneity (Hsieh et al., 2020; Koloskova et al., 2020). Additionally, through the proof of theorem 3 in Yu et al. (2019), it can be shown that for the non-convex objective function $f_{i}(\mathbf{x})$, DmSGD will also converge to a neighborhood of local optimal, with the radius being influenced by the variance of the stochastic gradient $\sigma^{2}$ and the data heterogeneity $\zeta^{2}$. When the data exhibit greater heterogeneity, DmSGD tends to oscillate over a larger area surrounding the optimum, resulting in larger corresponding errors. Therefore, addressing the challenge of leveraging momentum to accelerate the decentralized algorithm while simultaneously reducing heterogeneity is an urgent problem that needs to be resolved.
3.2 From Exact-Diffusion to Proposed Method
The challenge of heterogeneity is inherent in distributed SGD. To ensure the stability of the convergence, we need to guarantee that the parameter at the optimal value point remains in the optimal position after one step of iteration. However, the heterogeneity of data commonly leads to the condition that $\nabla f_{i}(\mathbf{X}^{\star})\neq 0$ (where $\mathbf{X}^{\star}=\mathbf{1}{\mathbf{x}^{\star}}^{\top}$ and $\mathbf{x}^{\star}$ denotes the optimal point). Consequently, the algorithm will move away from the optimal value after one step gradient update, resulting in $\mathbf{X}^{(t+1)}=\mathbf{X}^{\star}-\alpha\nabla\mathbf{f}(\mathbf{X}^{\star})\neq\mathbf{X}^{\star}$. Consequently, it is essential to employ specific methods to correct the biases introduced by the inconsistency in the gradient. A class of bias-correction distributed SGD algorithms has been developed to address heterogeneity, summarized as the SUDA algorithm in Alghunaim and Yuan (2022). In this paper, we focus on a specific class of algorithms based on the ED/D2 framework, which can be summarized in the following three steps: first, perform local gradient descent represented by $\bm{\psi}_{i}^{(t+1)}=\mathbf{x}_{i}^{(t)}-\alpha\nabla{F}_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$, where we use $\bm{\psi}_{i}^{(t)}$ to denote the temporary variable. Second, correct the bias caused by the data heterogeneity as expressed by $\bm{\phi}_{i}^{(t+1)}=\bm{\psi}_{i}^{(t+1)}+\mathbf{x}_{i}^{(t)}-\bm{\psi}_{i}^{(t)}$. Finally, combine the corrected parameter $\bm{\phi}_{i}^{(t+1)}$ to update the local parameter with $\mathbf{x}_{i}^{(t+1)}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\bm{\phi}_{j}^{(t+1)}$, where $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$ represents the set of neighbors for agent $i$.
The main idea of ED/D2 is to utilize local historical information to correct the bias introduced by the data heterogeneity during the local gradient update step. Thus, ED/D2 will ultimately converge to a neighborhood of the optimum, influenced solely by the randomness inherent in the stochastic gradient. Generally, the iteration of ED/D2 can be expressed as follows:
$\mathbf{X}^{(t+2)}=$ $\mathbf{W}\left(2\mathbf{X}^{(t+1)}-\mathbf{X}^{(t)}-\alpha\nabla\mathbf{F}(\mathbf{X}^{(t+1)},\bm{\Xi}^{(t+1)})+\alpha\nabla\mathbf{F}(\mathbf{X}^{(t)},\bm{\Xi}^{(t)})\right).$
To incorporate the idea of momentum acceleration into ED/D2, we replace $\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$ with a weighted sum over the historical gradient, specifically $\mathbf{m}_{i}^{(t)}$ from DmSGD as defined in (3.2). This leads to the development of the Exact-Diffusion with Momentum (EDM) algorithm, which can be formally described in Algorithm 1. The iteration formula can be written as:
$\mathbf{X}^{(t+2)}=\mathbf{W}\left(2\mathbf{X}^{(t+1)}-\mathbf{X}^{(t)}-\alpha\mathbf{M}^{(t+1)}+\alpha\mathbf{M}^{(t)}\right).$ (3.4)
Algorithm 1 Exact Diffusion with Momentum (EDM)
Initialization: Generate the initial value $\mathbf{x}^{(0)}$ from data or randomly and let $\mathbf{x}_{i}^{(-1)}=\mathbf{x}_{i}^{(0)}=\mathbf{x}^{(0)}$ at each node $i$. Initialize the parameters $\alpha$ and the double stochastic matrix $\mathbf{W}$.
for $t=0\cdots T-1$ do
     Estimate the local stochastic gradient based on $\bm{\xi}_{i}^{(t)}$: $\mathbf{g}_{i}^{(t)}=\nabla F_{i}(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i}^{(t)})$;
     Momentum update: $\mathbf{v}_{i}^{(t)}=\beta\mathbf{m}_{i}^{(t-1)}+(1-\beta)\mathbf{g}_{i}^{(t)}$;
     Adapt: $\bm{\psi}_{i}^{(t+1)}=\mathbf{x}_{i}^{(t)}-\alpha\mathbf{m}^{(t)}$;
     Correct: $\bm{\phi}_{i}^{(t+1)}=\bm{\psi}_{i}^{(t+1)}+\mathbf{x}_{i}^{(t)}-\bm{\psi}_{i}^{(t)}$;
     Combine: $\mathbf{x}_{i}^{(t+1)}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\bm{\phi}_{j}^{(t+1)}$;
end for
Output: $\mathbf{x}_{i}^{(T)}$ for all agent $i\in\mathcal{V}$.
Assume that $\mathbf{X}^{(t)}$ is obtained by Algorithm 1, and $\bar{\mathbf{x}}^{(t)}=\frac{1}{n}\mathbf{X}^{(t)}\mathbf{1}_{n}$, we have $\bar{\mathbf{x}}^{(t+1)}=\bar{\mathbf{x}}^{(t)}-\alpha\bar{\mathbf{m}}^{(t)}$. This implies that although the parameters differ across agents, the EDM algorithm exhibits the same update pattern as standard momentum SGD in the average sense.
4 Convergence Analysis
4.1 Auxiliary Variables
To measure the convergence rate, we primarily focus on $f(\bar{\mathbf{x}}^{(t)})$ and its gradient $\nabla f(\bar{\mathbf{x}}^{(t)})$. We begin by defining an auxiliary sequence for $\bar{\mathbf{x}}^{(t)}$, which is commonly used in the analysis of momentum-based SGD methods. Let $\mathbf{z}^{(0)}=\bar{\mathbf{x}}^{(0)}$ and define
$\mathbf{z}^{(t)}=\frac{1}{1-\beta}\bar{\mathbf{x}}^{(t)}-\frac{\beta}{1-\beta}\bar{\mathbf{x}}^{(t-1)},\quad t\geq 1.$ (4.5)
The following two lemmas establish the connection between $\mathbf{z}^{(t)}$ and $\mathbf{X}^{(t)}$ under the non-convex cases and PL conditions, respectively.
Lemma 1.
Suppose that Assumptions 1-3 hold. We have the following inequality holds for $t\geq 0$,
$\mathbb{E}[f(\mathbf{z}^{(t+1)})]\leq$ $\mathbb{E}[f(\mathbf{z}^{(t)})]+\frac{\alpha^{2}L\beta}{2(1-\beta)}\mathbb{E}\|\bar{\mathbf{m}}^{(t-1)}\|^{2}+\frac{\alpha L^{2}}{2n}\mathbb{E}\left\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}$ (4.6)
$-\frac{\alpha}{2}\left(1-\frac{\beta\alpha L}{1-\beta}-\alpha L\right)\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}-\frac{\alpha}{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\right\|^{2}.$
Lemma 2.
Let $\tilde{f}(\mathbf{x})=\sum_{i=1}^{n}f_{i}(\mathbf{x})-f^{\star}$. Under Assumptions 1-4, we have
$\mathbb{E}[\tilde{f}(\mathbf{z}^{(t+1)})]\leq$ $(1-\alpha\mu)\mathbb{E}[\tilde{f}(\mathbf{z}^{(t)})]-\frac{\alpha(1-\alpha L)}{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}$ (4.7)
$+\frac{\alpha L^{2}}{2n}\mathbb{E}\left\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{\alpha^{3}L^{2}\beta^{2}}{2(1-\beta)^{2}}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t-1)}\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}.$
4.2 Pseudo Deviation Variance Decomposition
According to Lemma 1 and 2, our next step is to conduct an in-depth investigation of $\mathbb{E}\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. To simplify the notation, we introduce $\mathbf{P}_{\mathbf{I}}=\mathbf{I}-\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}$, thus $\|\mathbf{X}^{(t)}-\overline{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}=\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|_{\mathrm{F}}^{2}$.
The momentum updating structure complicates the exploitation of the independence of $\bm{\Xi}^{(t)}$ in the stochastic gradient. To address this challenge, we introduce the following variables: let $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$ and $\widetilde{\mathbf{M}}^{(t)}=(1-\beta)\sum_{j=0}^{k}\beta^{t-j}\nabla\mathbf{f}(\mathbf{X}^{(j)})$, $\widetilde{\mathbf{X}}^{(-1)}=\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$ and
$\widetilde{\mathbf{X}}^{(t+1)}=\mathbf{W}\left(2\widetilde{\mathbf{X}}^{(t)}-\widetilde{\mathbf{X}}^{(t-1)}-\alpha\widetilde{\mathbf{M}}^{(t)}+\alpha\widetilde{\mathbf{M}}^{(t-1)}\right).$ (4.8)
By applying Cauchy’s inequality, we obtain:
$\left\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\right\|_{\mathrm{F}}^{2}\leq 2\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\right\|_{\mathrm{F}}^{2}+2\left\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\right\|_{\mathrm{F}}^{2}.$ (4.9)
Next, we present the upper bound of the noise term.
Lemma 3.
Suppose that Assumptions 1 and 3 hold. Let $\mathbf{X}^{(t)}$ is the sequence of parameters obtained from Algorithm 1 and $\widetilde{\mathbf{X}}^{(t)}$ be defined as above. Then, we have:
$\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\right\|_{\mathrm{F}}^{2}\leq\frac{13\alpha^{2}\lambda^{2}n\sigma^{2}}{1-\lambda}.$ (4.10)
This lemma provides an upper bound for $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\|_{\mathrm{F}}^{2}$, which is independent of $\beta$. Thus, we should only focus on the pseudo-deterministic part $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}$. We call it pseudo-deterministic because it eliminates randomness in the gradients while still retaining randomness in $\mathbf{X}^{(t)}$.
4.3 Transformations and Consensus Inequalities
Building on the ideas presented in Alghunaim and Yuan (2022), we can observe that $\widetilde{\mathbf{X}}^{(t)}$ exhibits a recursive structure similar to that of the SUDA framework described in their work. Specifically, we define $\widetilde{\mathbf{X}}^{(0)}=\mathbf{X}^{(0)}$, $\widetilde{\mathbf{M}}^{(0)}=(1-\beta)\nabla\mathbf{f}(\mathbf{X}^{(0)})$, $\widetilde{\mathbf{M}}^{(-1)}=\mathbf{0}$, $\widetilde{\mathbf{Y}}^{(0)}=\mathbf{0}$. For all $t\geq 0$, we have:
$\widetilde{\mathbf{X}}^{(t+1)}$ $=\mathbf{W}(\widetilde{\mathbf{X}}^{(t)}-\alpha\widetilde{\mathbf{M}}^{(t)})-(\mathbf{I}-\mathbf{W})^{1/2}\widetilde{\mathbf{Y}}^{(t)},$ (4.11)
$\widetilde{\mathbf{Y}}^{(t+1)}$ $=\widetilde{\mathbf{Y}}^{(t)}+(\mathbf{I}-\mathbf{W})^{1/2}\widetilde{\mathbf{X}}^{(t+1)},$ (4.12)
$\widetilde{\mathbf{M}}^{(t+1)}$ $=\beta\widetilde{\mathbf{M}}^{(t)}+(1-\beta)\nabla\mathbf{f}(\widetilde{\mathbf{X}}^{(t+1)}).$ (4.13)
The following lemma presents a transformation of $\widetilde{\mathbf{X}}^{(t)}$, which is used to derive the upper bound of $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|^{2}$, and consequently for $\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\|^{2}$. This lemma establishes an upper bound on the data consensus deviation that occurs during the iteration, which is central to our analysis of the convergence properties of algorithms.
Lemma 4.
Let $\widetilde{\mathbf{N}}^{(t)}$ be any series that satisfies $\widetilde{\mathbf{N}}^{(0)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$ and $\widetilde{\mathbf{R}}^{(t)}=\widetilde{\mathbf{M}}^{(t)}-\widetilde{\mathbf{N}}^{(t)}$. Under the conditions specified in assumptions 1-3, there exists a series of matrices $\{\mathbf{E}^{(t)}\}_{t\geq 1}$, satisfies the following two inequality
$\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}$ $\leq\sqrt{\lambda}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}\lambda}{1-\sqrt{\lambda}}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{2\alpha^{2}\lambda}{(1-\sqrt{\lambda})(1-\lambda)}\left\|\mathbf{P}_{\mathbf{I}}(\widetilde{\mathbf{N}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t)})\right\|_{\mathrm{F}}^{2},$ (4.14)
$\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\right\|_{\mathrm{F}}^{2}\leq 8\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{26\alpha^{2}\lambda^{2}n\sigma^{2}}{1-\lambda}.$ (4.15)
Remark 5.
In this Lemma, we do not specify the exact form of $\widetilde{\mathbf{N}}^{(t)}$. Because taking different values for $\widetilde{\mathbf{N}}^{(t)}$ in different proofs can simplify our proof, even though all forms are intended to measure the difference between the current updates and $\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$. In the subsequent proofs, we will adopt $\widetilde{\mathbf{N}}^{(t)}=(1-\beta)\sum_{j=0}^{t}\beta^{t-j}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(j)})+\beta^{t+1}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$ for the non-convex case and $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$ for the PL condition.
4.4 Non-Convex
The following Theorem presents the convergence properties under non-convex conditions.
Theorem 5.
Under Assumptions 1-3, if the step size $\alpha$ satisfies $\alpha\leq\min\left\{\frac{1-\sqrt{\lambda}}{4L},\frac{1-\beta}{4L}\right\}$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\bigg{(}\frac{1}{4}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{t}))\right\|^{2}+\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{t}))\right\|^{2}\bigg{)}$
$\leq$ $\frac{2}{\alpha T}\left(f(\mathbf{x}^{(0)})-f^{\star}\right)+\frac{2\alpha L\sigma^{2}}{n}+\frac{52\alpha^{2}L^{2}\lambda^{2}\sigma^{2}}{1-\lambda}+\frac{8C_{0}\alpha^{2}L^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}T}.$
where $\zeta_{0}^{2}=\frac{1}{n}\left\|\mathbf{W}\left(\mathbf{I}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\right)\nabla\mathbf{f}(\mathbf{X}^{(0)})\right\|_{\mathrm{F}}$, $f^{\star}=\min_{x}f(x)$ and $C_{0}=\frac{24(1-\beta)^{2}}{1+\sqrt{\lambda}}+12\beta^{2}\lambda(1-\sqrt{\lambda})+\frac{2\beta^{3}\lambda}{1-\beta}$.
In practical optimization problems, we typically select $\beta\in[0.8,0.99)$, allowing us to treat $\beta$ as a fixed value. When $\beta=0$, the algorithm simplifies to the ED/D2 method. Theorem 5 aligns with the latest conclusions of the ED/D2 framework discussed in Alghunaim and Yuan (2022) with the corresponding bounds illustrated in Table 1. In contrast to their findings, we obtained a bound that is unaffected by the minimum eigenvalue of the communication matrix through a refined scaling of the initial values. Consequently, we achieve a tighter upper bound when $\beta=0$. This results in the momentum-based algorithm exhibiting a convergence rate that is at least comparable to the ED/D2 algorithm. This feature shows the strength of our algorithm.
Remark 6.
Unlike Quasi-Global and DecentLaM, which are also the momentum-based variants of distributed stochastic gradient descent, EDM can ultimately eliminate the effects of data heterogeneity. Compared with another kind of bias-correction algorithm (DSGT and DSGT-HB), EDM can eliminate data heterogeneity at a rate of $O(\alpha^{2}(1-\lambda)^{-2}T^{-1})$, while DSGT at a rate of $O(\alpha^{2}(1-\lambda)^{-3}T^{-1})$. This distinction will be demonstrated in our subsequent real-data analysis.
4.5 PL-Condition
Theorem 6.
Suppose Assumptions 1-4 hold, if the step size $\alpha$ in the EDM algorithm satisfies $\alpha\leq\min\{\frac{1-\sqrt{\lambda}}{10},\frac{1-\beta}{5}\}$, we have
$\mathbb{E}f(\bar{\mathbf{x}}^{(t)})\leq\left(9\tilde{f}(\mathbf{x}^{(0)})+\frac{D_{1}\alpha^{3}L^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}}\right)\rho_{1}^{t}+\frac{D_{2}\alpha^{2}L\zeta_{0}^{2}}{1-\lambda}\rho_{2}^{t}+\frac{6\alpha L\sigma^{2}}{n\mu}+\frac{169\alpha^{2}L^{2}\lambda^{2}\sigma^{2}}{\mu(1-\lambda)}.$
where $\tilde{f}(\mathbf{x})=f(\mathbf{x})-f^{\star}$, $\rho_{1}=1-\alpha\mu$, $\rho_{2}=1-\min\left\{\frac{1-\sqrt{\lambda}}{5},\frac{2(1-\beta)}{5}\right\}$, and
$D_{1}=\frac{50\lambda\beta^{2}}{3(1-\beta)}+\frac{320(1-\beta)^{2}}{3(1+\sqrt{\lambda})}+\frac{160}{3}\lambda\beta^{2}(1-\sqrt{\lambda}),$
$D_{2}=\frac{100}{3}\left(\frac{5\lambda\beta^{2}}{3(1-\beta)}+4(1-\beta)^{2}+2\lambda\beta^{2}(1-\lambda)\right).$
Similarly to the analysis in the non-convex case, in the PL condition case, when $\beta=0$, our algorithm can degenerate to ED/D2. Moreover, our algorithm inherits the property of ED/D2, enabling it to linearly eliminate data heterogeneity.
Remark 7.
We compare EDM to other momentum-based distributed algorithms under the PL condition. Both the convergence bounds of Quasi-Global and DecentLaM fail to eliminate the effect of heterogeneity. Currently, no work proves the complete convergence properties of the DSGT algorithm with momentum under the Polyak-Lojasiewicz (PL) condition, or even under the strong convexity condition. While Huang et al. (2024) demonstrates convergence properties when $\beta$ is sufficiently large, this result does not extend to the original algorithm (i.e. when $\beta=0$). In contrast, our findings impose no constraints on the choice of $\beta$ and encompass the conclusions of ED/D2. Furthermore, our approach can be generalized to establish convergence properties for momentum-based DSGT algorithms.
On the other hand, Huang et al. (2024) employs the technique of Loopless Chebyshev Acceleration (LCA) to reduce the communication gap in DSGT from $(1-\lambda)^{-1}$ to $(1-\lambda)^{-1/2}$, resulting in a better upper bound on convergence. However, this technique is not exclusive to DSGT; other distributed algorithms can also benefit from LCA to reduce the communication gap. But this aspect is not the focus of our article.
5 Conclusions
This paper focuses on the acceleration algorithm of decentralized stochastic gradient descent in sparse communication networks and under conditions of data heterogeneity. By incorporating the idea of momentum acceleration into the bias-correction decentralized stochastic gradient algorithm, we propose a new algorithm that achieves consensus more rapidly in the presence of heterogeneous data. We theoretically demonstrate that, under non-convex and PL conditions, the proposed algorithm exhibits at least comparable convergence properties to the original algorithm, effectively mitigating the impact of data heterogeneity.
However, the specific mechanisms of momentum acceleration remain unclear. Therefore, further investigation is necessary to establish a robust framework for distributed momentum acceleration and, more broadly, for centralized momentum SGD.
Appendix A Some Basic Properties
Lemma 7.
Under Assumption 3, the following inequality holds,
$\sum_{t=0}^{T-1}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t)}\right\|^{2}\leq\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{(1-\beta)\sigma^{2}T}{n}.$ (A.16)
Proof of Lemma 7:   According to the iterative formula for $\mathbf{M}^{(t)}$, we have $\bar{\mathbf{m}}^{(t)}=\beta\bar{\mathbf{m}}^{(t-1)}+(1-\beta)\nabla\overline{\mathbf{F}}(\mathbf{X}^{(t)},\bm{\Xi}^{(t)})$. Noting that $\bm{\Xi}^{(t)}$ is independent of $\mathcal{F}^{(t)}$, we can get
$\mathbb{E}\left[\left\|\bar{\mathbf{m}}^{(t)}\right\|^{2}\Big{|}\mathcal{F}^{(t)}\right]\leq$ $\beta\left\|\bar{\mathbf{m}}^{(t-1)}\right\|^{2}+(1-\beta)\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{(1-\beta)^{2}\sigma^{2}}{n}.$ (A.17)
which follows from the Cauchy-Schwarz inequality and Assumption 3. Taking full expectation on both sides and iterating, then summing up both sides from 00 to $T-1$, the lemma holds.
Appendix B Proof of Lemmas
Proof of Lemma 1:   According to the $L$-smoothness of $f_{i}$, we have, we have
$\mathbb{E}[f(\mathbf{z}^{(t+1)})|\mathcal{F}^{(t)}]\leq f(\mathbf{z}^{(t)})+\langle\nabla f(\mathbf{z}^{(t)}),-\alpha\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\rangle+\frac{\alpha^{2}L}{2}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}.$ (B.18)
$\langle\nabla f(\mathbf{z}^{(t)}),-\alpha\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\rangle$ (B.19)
$\leq$ $\frac{1-\beta}{2L\beta}\left\|\nabla f(\mathbf{z}^{{(t)}})-\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\right\|^{2}+\frac{\alpha^{2}L\beta}{2(1-\beta)}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}-\alpha\langle\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)}),\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\rangle.$
The smoothness of $f_{i}$ and (4.5) implies $\|\nabla f(\mathbf{z}^{(t)})-\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\|^{2}\leq\frac{\alpha^{2}L^{2}\beta^{2}}{(1-\beta)^{2}}\|\bar{\mathbf{m}}^{(t-1)}\|^{2}$. Using the fact that $2\langle\mathbf{a},\mathbf{b}\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a-b}\|^{2}$, we have
$\langle\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)}),\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\rangle\geq\frac{1}{2}\left\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\right\|^{2}+\frac{1}{2}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}-\frac{L^{2}}{2n}\left\|\overline{\mathbf{X}}^{(t)}-\mathbf{X}^{(t)}\right\|_{\mathrm{F}}^{2}.$ (B.20)
Substituting them into (B.18) confirms the validity of the lemma.
Proof of Lemma 2:   Using $2\langle\mathbf{a},\mathbf{b}\rangle=\left\|\mathbf{a}\right\|^{2}+\left\|\mathbf{b}\right\|^{2}-\left\|\mathbf{a}-\mathbf{b}\right\|^{2}$ in (B.18) and $\left\|\nabla f(\mathbf{z}^{(t)})\right\|^{2}\leq 2\tilde{f}(\mathbf{z}^{(t)})$, we have
$\mathbb{E}[\tilde{f}(\mathbf{z}^{(t+1)})|\mathcal{F}^{(t)}]$
$\leq$ $\tilde{f}(\mathbf{z}^{(t)})-\frac{\alpha}{2}\left\|\nabla f(\mathbf{z}^{(t)})\right\|^{2}-\frac{\alpha(1-\alpha L)}{2}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha}{2}\left\|\nabla f(\mathbf{z}^{(t)})-\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}$
$\leq$ $(1-\alpha\mu)\tilde{f}(\mathbf{z}^{(t)})-\frac{\alpha(1-\alpha L)}{2}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha L^{2}}{2n}\sum_{i=1}^{n}\left\|\mathbf{z}^{(t)}-\mathbf{x}_{i}^{(t)}\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}$
$\leq$ $(1-\alpha\mu)\tilde{f}(\mathbf{z}^{(t)})-\frac{\alpha(1-\alpha L)}{2}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha L^{2}}{2n}\left\|\mathbf{P}_{\mathbf{I}}\mathbf{X}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{\alpha^{3}L^{2}\beta^{2}}{2(1-\beta)^{2}}\left\|\bar{\mathbf{m}}^{(t-1)}\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n}.$
Taking full expectation on both sides yields the desired result.
Proof of Lemma 3:   Let $\frac{1}{\alpha}\mathbf{U}^{\top}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})=(\mathbf{a}_{1}^{(t)},\cdots,\mathbf{a}_{n}^{(t)})^{\top}$ and $\bm{\Lambda}\mathbf{U}^{\top}(\widetilde{\mathbf{M}}^{(t)}-\mathbf{M}^{(t)})=(\mathbf{b}_{1}^{(t)},\cdots,\mathbf{b}_{n}^{(t)})^{\top}$. Taking the difference between (3.4) and (4.8), and multiplying both sides by $\mathbf{U}^{\top}$ we have
$\mathbf{a}_{i}^{(t+1)}=\lambda_{i}\left(2\mathbf{a}_{i}^{(t)}-\mathbf{a}_{i}^{(t-1)}\right)+\mathbf{b}_{i}^{(t)}-\mathbf{b}_{i}^{(t-1)}.$
When $\lambda_{i}=0$, we have $\mathbf{a}_{i}=0$. When $\lambda_{i}>0$, define $u_{i}=\sqrt{\lambda_{i}}\mathrm{e}^{j\theta}$ and $v_{i}=\sqrt{\lambda_{i}}\mathrm{e}^{-j\theta}$, where $\theta_{i}=\arccos(\sqrt{\lambda_{i}})$ and $j=\sqrt{-1}$. Let $c_{i}^{(t)}=\frac{u_{i}^{t}-v_{i}^{t}}{u_{i}-v_{i}}$ and $\delta^{(t)}=\mathbf{U}^{\top}(\nabla\mathbf{F}(\mathbf{X}^{(t)},\bm{\Xi}^{(t)})-\nabla\mathbf{f}(\mathbf{X}^{(t)}))=(\bm{\delta}_{1}^{(t)},\cdots,\bm{\delta}_{n}^{(t)})^{\top}$. Then
$\mathbf{a}_{i}^{(t)}=$ $\sum_{s=0}^{t-1}\left(\mathbf{b}_{i}^{(s)}-\mathbf{b}_{i}^{(s-1)}\right)\frac{u_{i}^{t-s}-v_{i}^{t-s}}{u_{i}-v_{i}}=\lambda_{i}(1-\beta)\sum_{r=0}^{t-1}\sum_{s=r}^{t-1}(c_{i}^{(t-s)}-c_{i}^{(t-s-1)})\beta^{s-r}\bm{\delta}_{i}^{(r)}$
$=$ $\frac{\lambda_{i}(1-\beta)}{u_{i}-v_{i}}\sum_{r=0}^{t-1}\bigg{(}\frac{(u_{i}-1)(u_{i}^{t-r}-\beta^{t-r})}{u_{i}-\beta}-\frac{(v_{i}-1)(v_{i}^{t-r}-\beta^{t-r})}{v_{i}-\beta}\bigg{)}\bm{\delta}_{i}^{(r)}$
$=$ $\underbrace{\lambda_{i}(1-\beta)\bm{\delta}_{i}^{(t-1)}}_{\Delta^{(t)}_{i,1}}+\underbrace{\frac{\lambda_{i}(1-\beta)}{u_{i}-v_{i}}\sum_{r=0}^{t-2}\bigg{(}\frac{(u_{i}-1)u_{i}^{t-r}}{u_{i}-\beta}-\frac{(v_{i}-1)v_{i}^{t-r}}{v_{i}-\beta}\bigg{)}\bm{\delta}_{i}^{(r)}}_{\Delta^{(t)}_{i,2}}+\underbrace{\frac{\lambda_{i}(1-\beta)^{2}}{\lambda_{i}-2\lambda_{i}\beta+\beta^{2}}\sum_{r=0}^{t-2}\beta^{t-r}\bm{\delta}_{i}^{(r)}}_{\Delta^{(t)}_{i,3}},$
where the last inequality is because of $\frac{u_{i}-1}{u_{i}-\beta}-\frac{v_{i}-1}{v_{i}-\beta}=\frac{(1-\beta)(u_{i}-v_{i})}{(u_{i}-\beta)(v_{i}-\beta)}=\frac{(1-\beta)(u_{i}-v_{i})}{\lambda_{i}-2\lambda_{i}\beta+\beta^{2}}$. Since $u_{i}\neq\beta$ and $v_{i}\neq\beta$, the sum of above series is reasonable. Applying the Cauchy inequality, it follows that
$\left\|\mathbf{a}_{i}^{(t)}\right\|^{2}\leq$ $\left\|\Delta^{(t)}_{i,1}\right\|^{2}+2\left\|\Delta^{(t)}_{i,2}\right\|^{2}+2\left\|\Delta^{(t)}_{i,3}\right\|^{2}.$ (B.21)
Since $\bm{\Xi}^{(0)},\cdots,\bm{\Xi}^{(t)}$ are independent, $\mathcal{F}^{(t)}$ is a martingale. We can check that, for any series $\{d_{r}\}_{r=0}^{\infty}$, we have
$\mathbb{E}\bigg{\|}\sum_{r=0}^{t}d_{r}\bm{\delta}^{(r)}\bigg{\|}^{2}=$ $\sum_{r=0}^{t}d_{r}^{2}\mathbb{E}\|\bm{\delta}^{(r)}\|^{2}.$ (B.22)
The inequality below will be frequently used in the following proof,
$\sum_{i=2}^{n}\mathbb{E}\|\bm{\delta}_{i}^{(t)}\|^{2}\leq n\sigma^{2}.$ (B.23)
For $\|\Delta^{(t)}_{i,1}\|^{2}$, applying (B.23), we have $\sum_{i=2}^{n}\mathbb{E}\|\Delta^{(t)}_{i,1}\|^{2}\leq\lambda^{2}(1-\beta)^{2}n\sigma^{2}$.
Applying (B.22) to $\mathbb{E}\|\Delta^{(t)}_{i,2}\|^{2}$, we have
$\mathbb{E}\left\|\Delta^{(t)}_{i,2}\right\|^{2}=$ $\lambda_{i}^{2}\sum_{r=0}^{t-2}\eta_{i}^{(t-r)}\mathbb{E}\|\bm{\delta}_{i}^{(r)}\|^{2},$ (B.24)
where $\eta^{(t-r)}$ is defined as follows. Noting that $u_{i}$ is the conjugate of $v_{i}$, we have
$\eta_{i}^{(t-r)}:=\left(\frac{1-\beta}{u_{i}-v_{i}}\left(\frac{(u_{i}-1)u_{i}^{t-r}}{u_{i}-\beta}-\frac{(v_{i}-1)v_{i}^{t-r}}{v_{i}-\beta}\right)\right)^{2}$ (B.25)
$\leq$ $\frac{4(1-\beta)^{2}}{\lambda_{i}-\lambda_{i}^{2}}\frac{|(u_{i}-1)(v_{i}-\beta)|^{2}\cdot|u_{i}^{t-r}|^{2}}{(\lambda_{i}-2\lambda_{i}\beta+\beta^{2})^{2}}=4\lambda_{i}^{t-r-1}(1-\lambda_{i})h_{1}^{2}(\beta,\lambda_{i})+4\lambda_{i}^{t-r}h_{2}^{2}(\beta,\lambda_{i}),$
where $h_{1}(\beta,\lambda_{i})=\frac{\beta(1-\beta)}{\lambda_{i}-2\lambda_{i}\beta+\beta^{2}}$ and $h_{2}(\beta,\lambda_{i})=\frac{(1-\beta)^{2}}{\lambda_{i}-2\lambda_{i}\beta+\beta^{2}}$. After some simple analysis, we can get for all $\lambda\in(0,1)$ and $\beta\in[0,1)$, $h_{1}(\beta,\lambda_{i})\leq\frac{1}{2\sqrt{\lambda_{i}(1-\lambda_{i})}}$ and $h_{2}(\beta,\lambda_{i})\leq\frac{1}{\lambda_{i}}$. Then we can get $\eta_{i}^{(t)}\leq 5\lambda_{i}^{t-r-2}$. Substituting it into (B.24) and summing the inequalities over $i$ from $2$ to $n$, we can get $\sum_{i=2}^{n}\mathbb{E}\left\|\Delta^{(t)}_{i,2}\right\|^{2}\leq\frac{5\lambda^{2}n\sigma^{2}}{1-\lambda}$.
Finally, we consider the bound of $\mathbb{E}[\|\Delta^{(t)}_{i,3}\|^{2}]$. Using (B.22) again, we have
$\sum_{i=2}^{n}\mathbb{E}\left[\left\|\Delta^{(t)}_{i,3}\right\|^{2}\right]\leq$ $\frac{(1-\beta)}{1-\lambda_{i}}\sum_{i=2}^{n}\lambda_{i}^{2}h_{3}^{2}(\beta,\lambda_{i})\mathbb{E}\|\bm{\delta}_{i}^{(r)}\|^{2}\leq\frac{\lambda^{2}(1-\beta)n\sigma^{2}}{1-\lambda}.$
where $h_{3}(\beta,\lambda)=\frac{\sqrt{1-\lambda}(1-\beta)\beta^{2}}{\lambda-2\lambda\beta+\beta^{2}}$ and the last inequality is because of $h_{3}(\beta,\lambda)\leq 1$.
Summing (B.21) from $2$ to $n$, combining the upper bound of $\sum_{i=2}^{n}\mathbb{E}\|\Delta^{(t)}_{i,k}\|^{2}$,$k=1,2,3$, and noting that $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}(\mathbf{X}^{(t)}-\widetilde{\mathbf{X}}^{(t)})\|_{\mathrm{F}}^{2}=\alpha^{2}\sum_{i=2}^{n}\mathbb{E}\|\mathbf{a}_{i}^{(t)}\|^{2}$, we can get the desired result.
Proof of Lemma 4:   Define $\widetilde{\mathbf{S}}^{(t)}=(\mathbf{I}-\mathbf{W})^{1/2}\widetilde{\mathbf{Y}}^{(t-1)}+\alpha\mathbf{W}\tilde{m}^{(t)}$. Let $\widetilde{\mathbf{N}}^{(t)}$ be any series that satisfies $\widetilde{\mathbf{N}}^{(0)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$. For $t\geq 1$, we have
$\widetilde{\mathbf{X}}^{(t+1)}=$ $(2\mathbf{W}-\mathbf{I})\widetilde{\mathbf{X}}^{(t)}-\widetilde{\mathbf{S}}^{(t)}-\alpha\mathbf{W}\widetilde{\mathbf{R}}^{(t)},$ (B.26)
$\widetilde{\mathbf{S}}^{(t+1)}=$ $\widetilde{\mathbf{S}}^{(t)}+(\mathbf{I}-\mathbf{W})\widetilde{\mathbf{X}}^{(t)}+\alpha\mathbf{W}(\widetilde{\mathbf{N}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t)}).$ (B.27)
If we define $\widetilde{\mathbf{Y}}^{(-1)}=\mathbf{0}$, we can check that (B.26) and (B.27) also hold for $t=0$.
Let $\mathbf{p}_{i}^{(t)}=[\mathbf{U}\widetilde{\mathbf{X}}^{(t)}]_{i,\cdot}$ and $\mathbf{q}_{i}^{(t)}=[(\mathbf{I}-\hat{\bm{\Lambda}})^{-1/2}\hat{\mathbf{U}}^{\top}\mathbf{S}^{(t)}]_{i-1,\cdot}$, $i\geq 2$; $\mathbf{z}_{i}^{(t)}=[\mathbf{U}^{\top}\widetilde{\mathbf{N}}^{(t)}]_{i,\cdot}$ and $\mathbf{r}_{i}^{(t)}=[\mathbf{U}^{\top}\widetilde{\mathbf{R}}^{(t)}]_{i,\cdot}$, where $[\mathbf{A}]_{i,\cdot}$ denotes the column vector of the $i$th row of matrix $\mathbf{A}$. Multiplying both sides of (B.26) and (B.27) by $\mathbf{U}^{\top}$, and with some transformation, we can get
$\mathbf{p}_{i}^{(t+1)}$ $=(2\lambda_{i}-1)\mathbf{p}_{i}^{(t)}-\sqrt{1-\lambda_{i}}\mathbf{q}_{i}^{(t)}-\alpha\lambda_{i}\mathbf{r}_{i}^{(t)},$
$\mathbf{q}_{i}^{(t+1)}$ $=\mathbf{q}_{i}^{(t)}+\sqrt{1-\lambda_{i}}\mathbf{p}_{i}^{(t)}+\frac{\alpha\lambda_{i}}{\sqrt{1-\lambda_{i}}}(\mathbf{z}_{i}^{(t+1)}-\mathbf{z}_{i}^{(t)}).$
Let
$G_{i}=$ $\begin{bmatrix}2\lambda_{i}-1&-\sqrt{1-\lambda_{i}}\\ \sqrt{1-\lambda_{i}}&1\end{bmatrix}=V_{i}\Gamma_{i}V_{i}^{-1},\quad V_{i}=\begin{bmatrix}-1&-1\\ \sqrt{1-\lambda_{i}}+j\sqrt{\lambda_{i}}&\sqrt{1-\lambda_{i}}-j\sqrt{\lambda_{i}}\end{bmatrix},$
and $\Gamma_{i}=\mathrm{diag}\{\lambda_{i}+\sqrt{\lambda_{i}^{2}-\lambda_{i}},\lambda_{i}-\sqrt{\lambda_{i}^{2}-\lambda_{i}}\}$, where $\mathrm{diag}\{a_{1},\cdots,a_{n}\}$ represents the diagonal matrix with elements $a_{1},\cdots,a_{n}$ on the diagonal. Alghunaim and Yuan (2022) proved that, when $10<\lambda_{i}<$ 1, we have $\|\Gamma_{i}\|_{\mathrm{op}}\leq\sqrt{\lambda_{i}}$, $\left\|V_{i}^{-1}\right\|_{\mathrm{op}}^{2}\leq\frac{1}{\lambda_{i}}$ and $\left\|V_{i}\right\|_{\mathrm{op}}^{2}\leq 4$.
Let $\mathbf{E}^{(t)}_{i}=V_{i}^{-1}(\mathbf{p}_{i}^{(t)},\mathbf{q}_{i}^{(t)})^{\top}$, $i=2,\cdots,n$, and $\mathbf{E}^{(t)}=((\mathbf{E}^{(t)}_{2})^{\top},\cdots,(\mathbf{E}^{(t)}_{n})^{\top})^{\top}$, then
$\mathbf{E}_{i}^{(t+1)}=\Gamma_{i}\mathbf{E}_{i}^{(t)}+\alpha\lambda_{i}V_{i}^{-1}\begin{bmatrix}-\mathbf{r}_{i}^{(t)}\\ (\mathbf{z}_{i}^{(t+1)}-\mathbf{z}_{i}^{(t)})/\sqrt{1-\lambda_{i}}\end{bmatrix}.$ (B.28)
Using the fact that $\left\|\mathbf{a}-\mathbf{b}\right\|^{2}\leq\frac{1}{\sqrt{\lambda}}\left\|\mathbf{a}\right\|^{2}+\frac{1}{1-\sqrt{\lambda}}\left\|\mathbf{b}\right\|^{2}$, and noting that $\left\|\mathbf{A}\mathbf{x}\right\|^{2}\leq\left\|\mathbf{A}\right\|_{\mathrm{op}}^{2}\left\|\mathbf{x}\right\|^{2}$, we have
$\left\|\mathbf{E}_{i}^{(t+1)}\right\|^{2}\leq$ $\sqrt{\lambda}\left\|\mathbf{E}_{i}^{(t)}\right\|^{2}+\frac{\alpha^{2}\lambda}{1-\sqrt{\lambda}}\left\|\mathbf{r}_{i}^{(t)}\right\|^{2}+\frac{\alpha^{2}\lambda}{(1-\sqrt{\lambda})(1-\lambda)}\left\|\mathbf{z}_{i}^{(t+1)}-\mathbf{z}_{i}^{(t)}\right\|^{2}.$ (B.29)
The following two inequalities give the upper bounds of $\sum_{i=2}^{n}\|\mathbf{r}_{i}^{(t)}\|^{2}$ and $\sum_{i=2}^{n}\|\mathbf{z}_{i}^{(t+1)}-\mathbf{z}_{i}^{(t)}\|^{2}$.
$\sum_{i=2}^{n}\left\|\mathbf{r}_{i}^{(t)}\right\|^{2}=\left\|\hat{\mathbf{U}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}=\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2},$ (B.30)
$\sum_{i=2}^{n}\left\|\mathbf{z}_{i}^{(t+1)}-\mathbf{z}_{i}^{(t)}\right\|^{2}=\left\|\mathbf{P}_{\mathbf{I}}\left(\widetilde{\mathbf{N}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t)}\right)\right\|_{\mathrm{F}}^{2}.$ (B.31)
Summing both sides of the (B.29) according to the subscript from 2 to $n$, and applying (B.30) and B.31 leads to (4.14) holds. And for (4.15), we have $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{X}}^{(t)}\|_{\mathrm{F}}^{2}=\sum_{i=2}^{n}\|V_{i}\mathbf{E}_{i}^{(t)}\|^{2}\leq 4\|\mathbf{E}^{(t)}\|_{\mathrm{F}}^{2}.$ Substituting it into (9) and applying Lemma 3, we obtain the desired result. This completes the proof.
Appendix C Proof of Theorem 5
Proof of Theorem 5:   In this part we choose $\widetilde{\mathbf{N}}^{(t)}=(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(i)})+\beta^{t+1}\nabla\mathbf{f}(\overline{\mathbf{X}}^{(0)})$. Noting that $\sum_{r=0}^{t}(1-\beta)\beta^{t-r}+\beta^{t+1}=1$ and the convexity of $\left\|\cdot\right\|_{\mathrm{F}}^{2}$, we have the following inequality holds from the Jenson’s inequality
$\left\|\mathbf{P}_{\mathbf{I}}\left(\widetilde{\mathbf{N}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t)}\right)\right\|_{\mathrm{F}}^{2}=\bigg{\|}\sum_{r=0}^{t}(1-\beta)\beta^{t-r}\mathbf{P}_{\mathbf{I}}\left(\nabla\mathbf{f}(\overline{\mathbf{X}}^{(r+1)})-\nabla\mathbf{f}(\overline{\mathbf{X}}^{(r)})\right)\bigg{\|}_{\mathrm{F}}^{2}$ (C.32)
$\leq$ $\sum_{r=0}^{t}(1-\beta)\beta^{t-r}\left\|\mathbf{P}_{\mathbf{I}}\left(\nabla\mathbf{f}(\overline{\mathbf{X}}^{(r+1)})-\nabla\mathbf{f}(\overline{\mathbf{X}}^{(r)})\right)\right\|_{\mathrm{F}}^{2}\leq n\alpha^{2}L^{2}\sum_{r=0}^{t}(1-\beta)\beta^{t-r}\left\|\bar{\mathbf{m}}^{(r)}\right\|^{2},$
Let $\bm{\psi}^{(t)}=\sum_{r=0}^{t}(1-\beta)\beta^{t-r}\left\|\bar{\mathbf{m}}^{(r)}\right\|^{2}$. Taking (C.32) into (4.14), we can get
$\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}\leq\sqrt{\lambda}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\lambda}{1-\sqrt{\lambda}}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{n\alpha^{4}L^{2}\lambda}{(1-\sqrt{\lambda})(1-\lambda)}\bm{\psi}^{(t)}.$ (C.33)
Next, we consider the iteration bound for $\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\|_{\mathrm{F}}^{2}$,
$\widetilde{\mathbf{R}}^{(t+1)}=\widetilde{\mathbf{M}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t+1)}=\beta\widetilde{\mathbf{R}}^{(t)}+(1-\beta)\left(\nabla\mathbf{f}(\mathbf{X}^{(t+1)})-\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})\right).$
Applying the $L$-smoothness of $\nabla f_{i}(\mathbf{x})$ and Lemma 4, we have
$\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t+1)}\right\|_{\mathrm{F}}^{2}\leq\beta\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+(1-\beta)\mathbb{E}\left\|\nabla\mathbf{f}(\mathbf{X}^{(t+1)})-\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t+1)})\right\|_{\mathrm{F}}^{2}$ (C.34)
$\leq$ $\beta\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+(1-\beta)L^{2}\mathbb{E}\left\|\mathbf{X}^{(t+1)}-\overline{\mathbf{X}}^{(t+1)}\right\|_{\mathrm{F}}^{2}$
$\leq$ $\beta\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+8(1-\beta)L^{2}\mathbb{E}\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}+\frac{26n\alpha^{2}L^{2}\lambda^{2}(1-\beta)\sigma^{2}}{1-\lambda}.$
Let
$\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}=\begin{bmatrix}\sqrt{\lambda}&\frac{\alpha^{2}\lambda}{1-\sqrt{\lambda}}\\ 8(1-\beta)L^{2}&\beta\end{bmatrix},\quad\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}=\begin{bmatrix}\frac{n\alpha^{4}L^{2}\lambda}{(1-\sqrt{\lambda})(1-\lambda)}\\ 0\end{bmatrix},\quad\tilde{\sigma}^{2}=\frac{26n\alpha^{2}L^{2}\lambda^{2}(1-\beta)\sigma^{2}}{1-\lambda}.$
Let $g(x)=\det(\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}-x\mathbf{I})=(x-\sqrt{\lambda})(x-\beta)-\frac{8\alpha^{2}L^{2}\lambda(1-\beta)}{1-\sqrt{\lambda}}$. Let $\lambda_{\mathbf{G},i},i=1,2$, be the solution of $g(x)=0$. When $\alpha\leq\frac{1-\sqrt{\lambda}}{4L}$, we have
$g(1)=$ $(1-\sqrt{\lambda})(1-\beta)-\frac{8\alpha^{2}L^{2}\lambda(1-\beta)}{1-\sqrt{\lambda}}>\frac{(1-\sqrt{\lambda})(1-\beta)}{2}.$
We have $\sqrt{\lambda}+\beta-1<\lambda_{\mathbf{G},1}<\lambda_{\mathbf{G},2}<1$, which implies $\rho(\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}})<1$. Then we have
$\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t+1)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}\leq$ $\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}+\mathbb{E}\bm{\psi}^{(t)}\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}+\begin{bmatrix}0\\ \tilde{\sigma}^{2}\end{bmatrix}$
$\leq$ $\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}^{(t)}\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}+\sum_{s=1}^{t}\mathbb{E}\bm{\psi}^{(s)}\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}^{t-s}\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}+\sum_{s=1}^{t}\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}}^{t-s}\begin{bmatrix}0\\ \tilde{\sigma}^{2}\end{bmatrix}.$
Then
$\sum_{k=1}^{T}\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}\leq$ $(\mathbf{I}-\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}})^{-1}\left(\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}+\sum_{t=1}^{T-1}\mathbb{E}\bm{\psi}^{(t)}\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}+\begin{bmatrix}0\\ (T-1)\tilde{\sigma}^{2}\end{bmatrix}\right).$
Noting that $\det(\mathbf{I}-\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}})=g(1)\geq\frac{(1-\sqrt{\lambda})(1-\beta)}{2}$, we have
$\sum_{t=1}^{T}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}\leq\frac{1}{\det(\mathbf{I}-\mathbf{G}_{\mathbf{E},\widetilde{\mathbf{R}}})}\begin{bmatrix}1-\beta&\frac{\alpha^{2}\lambda}{1-\sqrt{\lambda}}\end{bmatrix}\left(\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}+\sum_{t=1}^{T-1}\mathbb{E}\bm{\psi}^{(t)}\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}+\begin{bmatrix}0\\ (T-1)\tilde{\sigma}^{2}\end{bmatrix}\right)$ (C.35)
$\leq$ $\begin{bmatrix}\frac{2}{1-\sqrt{\lambda}}&\frac{2\alpha^{2}\lambda}{(1-\sqrt{\lambda})^{2}(1-\beta)}\end{bmatrix}\left(\begin{bmatrix}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}\\ \mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}\end{bmatrix}+\sum_{t=1}^{T-1}\mathbb{E}\bm{\psi}^{(t)}\mathbf{b}_{\mathbf{E},\widetilde{\mathbf{R}}}+\begin{bmatrix}0\\ (T-1)\tilde{\sigma}^{2}\end{bmatrix}\right)$
$\leq$ $\frac{2}{1-\sqrt{\lambda}}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}+\frac{2n\alpha^{4}L^{2}\lambda}{(1-\sqrt{\lambda})^{2}(1-\lambda)}\sum_{k=1}^{T-1}\mathbb{E}\bm{\psi}^{(k)}+\frac{2\alpha^{2}\lambda}{(1-\sqrt{\lambda})^{2}(1-\beta)}\left(\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}+(T-1)\tilde{\sigma}^{2}\right).$
Since $\widetilde{\mathbf{R}}^{(0)}=-\beta\nabla\mathbf{f}(\mathbf{X}^{(0)})$, we have $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(0)}\|_{\mathrm{F}}^{2}\leq n\lambda^{2}\beta^{2}\zeta_{0}^{2}$. Taking it into (C.34), where we choose $k=0$ we can get
$\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(1)}\right\|_{\mathrm{F}}^{2}\leq n\beta^{3}\zeta_{0}^{2}+8(1-\beta)L^{2}\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}+\tilde{\sigma}^{2}.$ (C.36)
Next, we consider the bound of $\mathbb{E}\|\mathbf{E}^{(1)}\|_{\mathrm{F}}^{2}$. Take the norm of the iterative expression (B.28), and from $\left\|\mathbf{A}+\mathbf{B}\right\|_{\mathrm{F}}^{2}\leq 2\left\|\mathbf{A}\right\|_{\mathrm{F}}^{2}+2\left\|\mathbf{B}\right\|_{\mathrm{F}}^{2}$, we can get
$\left\|\mathbf{E}_{i}^{(1)}\right\|^{2}\leq$ $2\left\|\Gamma_{i}\right\|_{\mathrm{op}}^{2}\left\|\mathbf{E}_{i}^{(0)}\right\|^{2}+2\alpha^{2}\lambda_{i}^{2}\left\|V_{i}^{-1}\right\|_{\mathrm{op}}^{2}\left(\left\|\mathbf{r}_{i}^{(0)}\right\|^{2}+\frac{1}{1-\lambda_{i}}\left\|\mathbf{z}_{i}^{(1)}-\mathbf{z}_{i}^{(0)}\right\|^{2}\right)$ (C.37)
$\leq$ $2\lambda_{i}\left\|V_{i}^{-1}(\mathbf{p}_{i}^{(0)},\mathbf{q}_{i}^{(0)})^{\top}\right\|_{\mathrm{F}}^{2}+4\alpha^{2}\lambda\left\|\mathbf{r}_{i}^{(0)}\right\|^{2}+\frac{4\alpha^{2}\lambda}{1-\lambda}\left\|\mathbf{z}_{i}^{(1)}-\mathbf{z}_{i}^{(0)}\right\|^{2}$
$\leq$ $8\left\|\mathbf{p}_{i}^{(0)}\right\|^{2}+8\left\|\mathbf{q}_{i}^{(0)}\right\|^{2}+4\alpha^{2}\lambda\left\|\mathbf{r}_{i}^{(0)}\right\|^{2}+\frac{4\alpha^{2}\lambda}{1-\lambda}\left\|\mathbf{z}_{i}^{(1)}-\mathbf{z}_{i}^{(0)}\right\|^{2}.$
By the definition of $\mathbf{p}$, $\mathbf{q}$,$\mathbf{r}$ and $\mathbf{z}$, we have
$\sum_{i=2}^{n}\left\|\mathbf{p}_{i}^{(0)}\right\|^{2}=\left\|\widehat{\mathbf{U}}^{\top}\mathbf{X}^{(0)}\right\|^{2}=0,\quad\sum_{i=2}^{n}\left\|\mathbf{r}_{i}^{(0)}\right\|^{2}=\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(0)}\right\|_{\mathrm{F}}^{2}\leq n\beta^{2}\zeta_{0}^{2},$
$\sum_{i=2}^{n}\left\|\mathbf{q}_{i}^{(0)}\right\|^{2}\leq\frac{1}{1-\lambda}\left\|\widehat{\mathbf{U}}^{\top}\mathbf{s}^{(0)}\right\|^{2}\leq\frac{\alpha^{2}}{1-\lambda}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{N}}^{(0)}\right\|^{2}\leq\frac{n\alpha^{2}(1-\beta)^{2}\zeta_{0}^{2}}{1-\lambda}.$
And for the bound of $\sum_{i=2}^{n}\|\mathbf{z}_{i}^{(1)}-\mathbf{z}_{i}^{(0)}\|^{2}=\|\mathbf{P}_{\mathbf{I}}(\widetilde{\mathbf{N}}^{(1)}-\widetilde{\mathbf{N}}^{(0)})\|_{\mathrm{F}}^{2}$, we use (C.32). Then we have
$\mathbb{E}\left\|\mathbf{E}^{(1)}\right\|_{\mathrm{F}}^{2}=\sum_{i=2}^{n}\left\|\mathbf{E}_{i}^{(1)}\right\|^{2}\leq\frac{8n\alpha^{2}(1-\beta)^{2}\zeta_{0}^{2}}{1-\lambda}+4n\alpha^{2}\lambda\beta^{2}\zeta_{0}^{2}+\frac{4n\alpha^{4}L^{2}\lambda\bm{\psi}^{(0)}}{1-\lambda}.$ (C.38)
Substituting (C.36) and (C.38) into (C.35), we can get
$\sum_{t=1}^{T}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}\leq$ $\frac{n\alpha^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}}C_{0}+\frac{2n\alpha^{4}L^{2}\lambda}{(1-\sqrt{\lambda})^{2}(1-\lambda)}\sum_{t=0}^{T-1}\mathbb{E}\bm{\psi}^{(t)}+\frac{52Tn\alpha^{4}L^{2}\lambda^{3}\sigma^{2}}{(1-\sqrt{\lambda})^{2}(1-\lambda)},$ (C.39)
where $C_{0}=\frac{24(1-\beta)^{2}}{1+\sqrt{\lambda}}+12\beta^{2}\lambda(1-\sqrt{\lambda})+\frac{2\beta^{3}\lambda}{1-\beta}$.
Next, we consider the upper bound of $\sum_{k=0}^{T-1}\mathbb{E}\bm{\psi}^{(k)}$. Based on the definition of $\bm{\psi}$ and utilizing Lemma (A.16), we can obtain
$\sum_{k=0}^{T-1}\mathbb{E}\bm{\psi}^{(k)}=$ $\sum_{k=0}^{T-1}\sum_{s=0}^{k}(1-\beta)\beta^{s-k}\mathbb{E}\left\|\bar{\mathbf{m}}^{(k)}\right\|^{2}\leq\sum_{k=0}^{T-1}\mathbb{E}\left\|\bar{\mathbf{m}}^{(k)}\right\|^{2}\leq\sum_{k=0}^{T-1}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(k)})\right\|^{2}+\frac{(1-\beta)\sigma^{2}T}{n}.$ (C.40)
When $\alpha L\leq\frac{1-\beta}{4}$, we can rewrite (4.6) as
$\mathbb{E}[f(\mathbf{z}^{(t+1)})]\leq$ $\mathbb{E}[f(\mathbf{z}^{(t)})]+\frac{\alpha^{2}L\beta}{2(1-\beta)}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t-1)}\right\|^{2}-\frac{3\alpha}{8}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}$ (C.41)
$+\frac{\alpha^{2}L\sigma^{2}}{2n}-\frac{\alpha}{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\right\|^{2}+\frac{\alpha L^{2}}{2n}\mathbb{E}\left\|\overline{\mathbf{X}}^{(t)}-\mathbf{X}^{(t)}\right\|_{\mathrm{F}}^{2}.$
Combining (A.16), (C.39), (C.40),(C.41) and (4.14), we have
$\frac{\alpha}{2}\sum_{t=0}^{T-1}\left(\frac{1}{4}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{t})\right\|^{2}+\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{t})\right\|^{2}\right)\leq f(\mathbf{x}^{(0)})-\mathbb{E}[f(\mathbf{z}^{(T)})]+\frac{\alpha^{2}L\beta}{2(1-\beta)}\sum_{t=0}^{T-1}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t-1)}\right\|^{2}$
$-\frac{\alpha}{4}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}T}{2n}+\frac{4\alpha L^{2}}{n}\sum_{t=1}^{T-1}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{13T\alpha^{3}L^{2}\lambda^{2}\sigma^{2}}{1-\lambda}$
$\leq$ $f(\mathbf{x}^{(0)})-\mathbb{E}[f(\mathbf{z}^{(T)})]+\frac{\alpha^{2}L\beta\sigma^{2}T}{2n}+\frac{\alpha^{2}L\beta}{2(1-\beta)}\sum_{t=0}^{T-2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}-\frac{\alpha}{4}\sum_{t=0}^{T-1}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}T}{2n}$
$+C_{0}\frac{4\alpha^{3}L^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}}+\frac{8\alpha^{5}L^{4}}{(1-\sqrt{\lambda})^{2}(1-\lambda)}\sum_{k=0}^{T-1}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(k)})\right\|^{2}$
$+\frac{8T\alpha^{5}L^{4}\lambda(1-\beta)\sigma^{2}}{n(1-\sqrt{\lambda})^{2}(1-\lambda)}+\frac{208T\alpha^{5}L^{4}\lambda^{3}(1-\beta)\sigma^{2}}{(1-\sqrt{\lambda})^{2}(1-\lambda)}+\frac{13T\alpha^{3}L^{2}\lambda^{2}\sigma^{2}}{1-\lambda}$
$\leq$ $f(\mathbf{x}^{(0)})-\mathbb{E}[f(\mathbf{z}^{(T)})]+\frac{\alpha^{2}L\sigma^{2}T}{n}+\frac{4C_{0}\alpha^{3}L^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}}+\frac{26T\alpha^{3}L^{2}\lambda^{2}\sigma^{2}}{1-\lambda}.$
Dividing both sides of the equation by $\alpha T/2$, we finally obtain the desired result.
Appendix D Proof of Theorem 6
Lemma 8.
For any $L$-smooth function $f$, we have
$\left\|\nabla f(\mathbf{x})\right\|^{2}\leq 2L(f(\mathbf{x})-f^{\star}).$
Proof of Lemma 8:   See Polyak (1987).
Next, we prove Theorem 6.
Proof of Theorem 6:   In this part, we choose $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\bar{\mathbf{X}}^{(t)})$. Let $\{\mathbf{E}^{(t)}\}_{t\geq 1}$ be the series generated from the proof of lemma (10) with $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\bar{\mathbf{X}}^{(t)})$. Then form assumption 2, we have
$\left\|\mathbf{P}_{\mathbf{I}}(\widetilde{\mathbf{N}}^{(t+1)}-\widetilde{\mathbf{N}}^{(t)})\right\|_{\mathrm{F}}^{2}\leq n\alpha^{2}L^{2}\left\|\bar{\mathbf{m}}^{(t)}\right\|^{2}.$
Substitute it into (14) , we derive that
$\mathbb{E}\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}\leq\sqrt{\lambda}\mathbb{E}\left\|\mathbf{E}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{\alpha^{2}\lambda}{1-\sqrt{\lambda}}\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{n\alpha^{4}L^{2}\lambda}{(1-\sqrt{\lambda})(1-\lambda)}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t)}\right\|^{2}.$ (D.42)
Next we consider the iteration bound for $\mathbb{E}\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\|_{\mathrm{F}}$ under the newly defined $\tilde{\mathbf{m}}^{(t)}$. The iteration formula of $\widetilde{\mathbf{R}}^{(t)}$ can be written as follows
$\widetilde{\mathbf{R}}^{(t+1)}=\beta\widetilde{\mathbf{R}}^{(t)}+\beta\left(\nabla\mathbf{f}(\bar{\mathbf{X}}^{(t)})-\nabla\mathbf{f}(\bar{\mathbf{X}}^{(t+1)})\right)+(1-\beta)\left(\nabla\mathbf{f}(\mathbf{X}^{(t+1)})-\nabla\mathbf{f}(\bar{\mathbf{X}}^{(t+1)})\right).$
Using Jensen’s inequality (the weight of each item is $\beta,\frac{1-\beta}{2}$ and $\frac{1-\beta}{2}$), and applying (15), we have
$\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t+1)}\right\|_{\mathrm{F}}^{2}\leq\beta\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t)}\right\|_{\mathrm{F}}^{2}+\frac{2n\alpha^{2}L^{2}\beta^{2}}{1-\beta}\mathbb{E}\left\|\bar{\mathbf{m}}^{(t)}\right\|^{2}+16(1-\beta)L^{2}\mathbb{E}\left\|\mathbf{E}^{(t+1)}\right\|_{\mathrm{F}}^{2}+\frac{52n\alpha^{2}L^{2}\lambda^{2}(1-\beta)\sigma^{2}}{1-\lambda}.$ (D.43)
For $t\geq 1$, denote $h^{(t)}_{1}=\mathbb{E}[\tilde{f}(\mathbf{z}^{(t)})]$, $h^{(t)}_{2}=\mathbb{E}\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(t-1)}\|_{\mathrm{F}}^{2}$, $h^{(t)}_{3}=\mathbb{E}\|\mathbf{E}^{(t)}\|_{\mathrm{F}}^{2}$, and $h^{(t)}_{4}=\mathbb{E}\|\bar{\mathbf{m}}^{(t-1)}\|^{2}$. Taking them into (7), and applying (15), we have
$h_{1}^{(t+1)}\leq$ $(1-\alpha\mu)h_{1}^{(t)}+\frac{4\alpha L^{2}}{n}h_{3}^{(t)}+\frac{\alpha^{3}L^{2}\beta^{2}}{2(1-\beta)^{2}}h_{4}^{(t)}$
$-\frac{\alpha(1-\alpha L)}{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{13\alpha^{3}L^{2}\lambda^{2}\sigma^{2}}{1-\lambda}+\frac{\alpha^{2}L\sigma^{2}}{2n}.$
From (D.43) and (A.17), we have
$h^{(t+1)}_{2}\leq\beta h^{(t)}_{2}+16(1-\beta)L^{2}h_{3}^{(t)}+\frac{2n\alpha^{2}L^{2}\beta^{2}}{1-\beta}h_{4}^{(t)}+\left(\frac{10(1-\beta)}{1-\lambda}+3\right)4n\lambda^{2}\alpha^{2}L^{2}\sigma^{2},$
$h_{4}^{(t+1)}\leq\beta h_{4}^{(t)}+(1-\beta)\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(t)})\right\|^{2}+\frac{(1-\beta)^{2}\sigma^{2}}{n}.$ (D.44)
By substituting the inequality into (D.42). Under the condition $\alpha L\leq\min\{\frac{1-\sqrt{\lambda}}{10},\frac{1-\beta}{5}\}$, we have the relation $\sqrt{\lambda}+\frac{16\alpha^{2}L^{2}\lambda(1-\beta)}{1-\sqrt{\lambda}}\leq\frac{8+17\sqrt{\lambda}}{25}$. This leads to
$h_{3}^{(t+1)}\leq$ $\frac{8+17\sqrt{\lambda}}{25}h_{3}^{(t)}+\frac{\alpha^{2}\lambda\beta}{1-\sqrt{\lambda}}h_{2}^{(t)}+\frac{n\alpha^{3}L\lambda\beta}{2(1-\sqrt{\lambda})}h_{4}^{(t)}$ (D.45)
$+\frac{52n\alpha^{4}L^{2}\lambda^{3}(1-\beta)\sigma^{2}}{(1-\sqrt{\lambda})(1-\lambda)}+\frac{\alpha^{4}L^{2}\lambda(1-\beta)^{2}\sigma^{2}}{(1-\sqrt{\lambda})(1-\lambda)}+\frac{n\alpha^{4}L^{2}\lambda(1-\beta)}{(1-\sqrt{\lambda})(1-\lambda)}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{x}^{(t)})\right\|^{2}.$
Denote $\mathcal{L}^{(t)}=h_{1}^{(t)}+\frac{15\alpha^{3}L^{2}\lambda}{n(1-\sqrt{\lambda})^{2}(1-\beta)}h_{2}^{(t)}+\frac{12\alpha L^{2}}{n(1-\sqrt{\lambda})}h_{3}^{(t)}+\frac{2\alpha^{3}L^{2}}{(1-\sqrt{\lambda})(1-\beta)^{2}}h_{4}^{(t)}$. Following the proof of theorem 3.8 in Huang et al. (2024), we can get, when $t\geq 1$,
$\mathcal{L}^{(t+1)}\leq$ $(1-\alpha\mu)\mathcal{L}^{(t)}+\frac{7\alpha^{2}L\sigma^{2}}{10n}+\frac{676\alpha^{3}L^{2}\lambda^{2}\sigma^{2}}{25(1-\lambda)}.$ (D.46)
Next, we consider the condition when $t=0$, from (7) , $\widetilde{\mathbf{R}}^{(0)}=-\beta\nabla\mathbf{f}(\mathbf{X}^{(0)})$ and $\overline{M}^{(0)}=(1-\beta)\nabla\mathbf{F}(\mathbf{X}^{(0)},\bm{\Xi}^{(0)})$, we have
$h_{1}^{(1)}=\mathbb{E}\tilde{f}(\mathbf{z}^{(1)})\leq(1-\alpha\mu)\tilde{f}(\mathbf{x}^{(0)})-\frac{\alpha(1-\alpha)}{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(0)})\right\|^{2}+\frac{\alpha^{2}L\sigma^{2}}{2n},$ (D.47)
Noting that $\sum_{i=2}^{n}\|\mathbf{z}_{i}^{(1)}-\mathbf{z}_{i}^{(0)}\|^{2}=\|\mathbf{P}_{\mathbf{I}}(\widetilde{\mathbf{N}}^{(1)}-\widetilde{\mathbf{N}}^{(0)})\|_{\mathrm{F}}^{2}=\|\mathbf{P}_{\mathbf{I}}(\nabla\mathbf{f}(\bar{\mathbf{X}}^{(1)})-\nabla\mathbf{f}(\bar{\mathbf{X}}^{(0)}))\|_{\mathrm{F}}^{2}$, substituting it into (C.37), we can get the upper bound of $h_{3}^{(1)}$,
$h_{3}^{(1)}=\sum_{i=2}^{n}\left\|\mathbf{E}_{i}^{(1)}\right\|^{2}\leq\frac{8n\alpha^{2}(1-\beta)^{2}\zeta_{0}^{2}}{1-\lambda}+4n\alpha^{2}\lambda\beta^{2}\zeta_{0}^{2}+\frac{4\alpha^{4}L^{2}\lambda(1-\beta)^{2}\left(n\left\|\bar{\mathbf{f}}(\mathbf{X}^{(0)})\right\|^{2}+\sigma^{2}\right)}{1-\lambda}.$ (D.48)
And we can check that $h^{(1)}_{2}=\mathbb{E}\left\|\mathbf{P}_{\mathbf{I}}\widetilde{\mathbf{R}}^{(0)}\right\|_{\mathrm{F}}^{2}\leq n\beta^{2}\zeta_{0}^{2}$ and $h_{4}^{(1)}\leq(1-\beta)^{2}\mathbb{E}\left\|\nabla\bar{\mathbf{f}}(\mathbf{X}^{(0)})\right\|^{2}+\frac{(1-\beta)^{2}\sigma^{2}}{n}$.
Taking them into the definition of $\mathcal{L}^{(1)}$, we have
$\mathcal{L}^{(1)}\leq$ $(1-\alpha\mu)\left(\tilde{f}(\mathbf{x}^{(0)})+\frac{D_{1}\alpha^{3}L^{2}\zeta_{0}^{2}}{(1-\sqrt{\lambda})^{2}}\right)+\frac{7\alpha^{2}L\sigma^{2}}{10n},$ (D.49)
where $D_{1}=\frac{50\lambda\beta^{2}}{3(1-\beta)}+\frac{320(1-\beta)^{2}}{3(1+\sqrt{\lambda})}+\frac{160}{3}\lambda\beta^{2}(1-\sqrt{\lambda})$.
By repeatedly (D.46) and applying (D.49), we ultimately obtain that
$\mathcal{L}^{(t)}\leq(1-\alpha\mu)^{t}\left(\tilde{f}(\mathbf{x}^{(0)})+\frac{D_{1}\alpha^{3}L^{2}\zeta_{0}^{2}}{(1-\lambda)^{2}}\right)+\frac{7\alpha L\sigma^{2}}{10n\mu}+\frac{676\alpha^{2}L^{2}\lambda^{2}\sigma^{2}}{25\mu(1-\lambda)}.$ (D.50)
Analogous to the scaling done for $\mathcal{L}^{(t)}$, we can scale $\mathcal{H}^{(t)}$ and $h_{4}^{(t)}$ following the proof of Theorem 4.2 in Huang et al. (2024), where we choose $\mathcal{H}^{(t)}=\frac{5\alpha^{2}\lambda}{3(1-\sqrt{\lambda})(1-\beta)}h_{2}^{(t)}+h_{3}^{(t)}+\frac{2n\alpha^{2}\lambda}{3(1-\sqrt{\lambda})(1-\beta)}h_{4}^{(t)}$. For the condition when $t=0$, we apply (C.37) and $\widetilde{\mathbf{N}}^{(t)}=\nabla\mathbf{f}(\overline{\mathbf{X}}^{(t)})$ for a more detailed scaling. Finally, we can obtain that the desired result.
Appendix E Simulations and Real-Data Analysis
In this section, we demonstrate the advantages of our algorithm through simulations. We compare our algorithm with the following two classes of algorithms: (i) momentum-free bias-correction algorithm (DSGT and ED/D2), and (ii) momentum-based algorithm (DmSGT-HB, DecentLaM, and Quasi-Global). The first type of algorithm is used to show that our algorithm accelerates the convergence of the original one, and the second type of algorithm is used to show that our algorithm has stronger adaptability to data heterogeneity and network sparsity. In the following three parts, we will simulate the square loss function, the general strongly convex objective function, and the non-convex objective function, respectively.
For the double stochastic communication matrix $\mathbf{W}=(w_{ij})_{1\leq i,j\leq n}$, we model the sparsity of the communication network using a ring graph, where each agent is arranged in a circular formation and can only communicate with its two immediate neighbors. We define the weights as follows: $w_{i,i}=1/2,w_{i,i+1}=w_{i,i-1}=1/4$ and $w_{i,j}=0$ otherwise. In this context, we treat $-1$ as $n$ and $n+1$ as $1$ to account for the cyclic nature of the graph. It can be verified that the spectral gap of the above ring graph satisfies $1-\lambda=\mathcal{O}(1/n^{2})$. Consequently, as the number of agents increases, the sparsity of the network significantly increases.
E.1 Quadratic Loss
We first consider a simple linear regression problem to visualize the issues caused by network sparsity and data heterogeneity. The loss functions are defined as follows:
$f_{i}(\mathbf{x},\bm{\xi})=\frac{1}{2}\mathbb{E}_{\mathbf{y}_{i}}\|\mathbf{y}_{i}-\mathbf{A}_{i}\mathbf{x}\|^{2},$
where $i=1,\cdots,n$, $\mathbf{x}\in\mathbb{R}^{d},\mathbf{y}_{i}\in\mathbb{R}^{p}$, and $\mathbf{A}_{i}=(\mathbf{a}_{i1},\cdots,\mathbf{a}_{ip})^{\top}\in\mathbb{R}^{p\times d}$ represents the corresponding covariates at node $i$, with the covariates $\mathbf{a}_{i,j}$ being i.i.d. samples generated from $\mathcal{N}(0,\mathbf{I}_{d})$. Here, we choose $d=10$ and $p=20$. The local optimum parameter $\mathbf{x}_{i}^{\star}$ generated as follows: first, let $\mathbf{u}_{i}$ be i.i.d. samples drawn from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. The exact minimizer is given by $\mathbf{x}^{\star}=\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\right)^{-1}\left(\sum_{i}\mathbf{A}_{i}^{\top}\mathbf{A}_{i}\mathbf{u}_{i}\right)$. We then define $\mathbf{x}_{i}^{\star}=\mathbf{x}^{\star}+(\mathbf{u}_{i}-\mathbf{x}^{\star})/c$, allowing us to control the data heterogeneity $\zeta^{2}=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x}^{\star})\|^{2}$ through the parameter $c$. The responses $\mathbf{y}_{i}$ are generated as follows, $\mathbf{y}_{i}^{(t)}=\mathbf{A}_{i}\mathbf{x}_{i}^{\star}+\bm{\epsilon}_{i}^{(t)}$, where $\bm{\epsilon}_{i}^{(t)}$ are i.i.d. drawn from $\mathcal{N}(\mathbf{0}_{p},\sigma^{2}\mathbf{I}_{p})$.
With $n=32$ ($\lambda=0.99$) and $\sigma^{2}=0.05$, we conducted 20 experiments for various values of $\zeta^{2}$, and the results are presented in Figure 1.
Figure 1: Quadratic loss function. We choose $\sigma^{2}=0.05$, $\alpha=0.05$, $\lambda=0.99$ and use different $\zeta^{2}$ to control the heterogeneity of $f_{i}$.
The experimental results indicate that all algorithms can converge linearly to a neighborhood of optimal values. However, as data heterogeneity increases, other momentum-based algorithms (DmSGD, Quasi-Global, and DecentLaM) become trapped in heterogeneous regions. Although the DSGT-HB algorithm effectively eliminates heterogeneity errors, its convergence rate is impacted by the original DSGT algorithm, resulting in momentum acceleration failure during the gradient consensus phase. In contrast, the EDM method retains the beneficial properties of the ED/D2, and the introduction of momentum improves the convergence rate.
E.2 General Strongly Convex Loss
We consider the $\ell_{2}$-regularized logistic regression problem to investigate the algorithm’s convergence in a general strongly convex case. Let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ be the local parameter of node $i$, and $\bm{\xi}_{i}=\{(\mathbf{u}_{ij}^{\top},v_{ij})^{\top}\}_{j=1,\cdots,m}$, where $u_{ij}\in\mathbb{R}^{d}$ is the covariate and $v_{ij}\in\{-1,1\}$ is the response. Given $\mathbf{x}_{i}$ and $\mathbf{u}_{ij}$, the response $v_{ij}$ is generated such that $v_{ij}=1$ with probability $1/(1+\exp(-\mathbf{x}_{i}^{\top}\mathbf{u}_{ij}))$ and $v_{ij}=-1$ otherwise. The loss function at agent $i$ can be expressed as
$f_{i}(\mathbf{x}_{i})=\frac{1}{m}\sum_{j=1}^{m}\log[1+\exp(-v_{ij}\mathbf{x}_{i}^{\top}\mathbf{u}_{ij})]+\frac{\mu}{2}\left\|\mathbf{x}_{i}\right\|^{2}.$
It can be verified that this loss function is $\mu$-strongly convex.
[…]
We utilize the full batch gradient $\nabla F(\mathbf{x}_{i},\bm{\xi}_{i})$ at every iteration. However, since we are using full batch samples, this results in no randomness. Thus, we add an additional i.i.d. noise term $\mathbf{s}_{i}^{(t)}$ drawn from $\mathcal{N}(0,\sigma_{s}^{2}I_{d})$ to $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})$. And use $\nabla F(\mathbf{x}_{i}^{(t)},\bm{\xi}_{i})+\mathbf{s}_{i}^{(t)}$ as the stochastic gradient. This method is similar to the one used in Alghunaim and Yuan (2022) to control the gradient noise.
For our simulations, we set $d=20$ and $m=2000$, using the $\ell_{2}$-norm of the gradient $\|\nabla\bar{\mathbf{f}}(\overline{\mathbf{X}}^{(t)})\|_{2}^{2}$ as a measure of convergence. The simulation is conducted 20 times on a ring graph with $n=32$ nodes. We compare our method against DSGT, ED/D2, DSGT-HB, and DmSGD. The results are presented in Figure 2.
Figure 2: Logistic regression with $l_{2}$-regularization, random noise version. We choose $\sigma_{s}^{2}=0.01$, $\alpha=0.5$, $\lambda=0.99$, and use $\sigma_{h}^{2}$ to control the variance of $\mathbf{x}_{i}^{\star}$, which reflects the heterogeneity.
The simulation results demonstrate that our method enhances the performance of ED/D2, achieving convergence to a narrower region of global optimal values while experiencing minimal influence from data heterogeneity.
E.3 Non-Convex Loss
Finally, we consider the non-convex condition. We use VGG-11 as the backbone network to train the classification task of the CIFAR10 dataset consisting of $N=50,000$ images of 10 labels, such as cats, airplanes, etc. The photos are $32\times 32$ size with three color channels. The criterion is the cross entropy. To quantitatively analyze the heterogeneity of the data in this dataset, we introduce the Dirichlet distribution, which is widely used to model heterogeneous scenarios for classification problems (Yurochkin et al., 2019; Lin et al., 2021). We generate $\mathbf{p}_{k}=(p_{k1},\cdots,p_{kn})$ for $k=1,\cdots,10$ from $\mathrm{Dirichlet}(\phi,\cdots,\phi)$, allocating $p_{ki}$ proportion of samples with label $k$ to agent $i$. The Dirichlet parameter $\phi$ is used to control the weights assigned to different categories of data across devices, serving as a measure of data heterogeneity. Specifically, a smaller value of $\phi$ indicates greater heterogeneity among the samples allocated to different agents. Our simulation for this part is primarily based on the code provided in Vogels et al. (2021).
Figure 3: Classification of CIFAR10 by VGG11. Learning rate $\alpha=0.1$, heterogeneity parameter $\phi=1.0$ (heterogeneous).
Figure 4: Classification of CIFAR10 by VGG11. Learning rate $\alpha=0.1$, heterogeneity parameter $\phi=0.1$ (very heterogeneous).
We considered two cases with $\phi=1$ (heterogeneous) and $\phi=0.1$ (highly heterogeneous). For $\phi=1$, the learning rate is reduced to 10% of its original value during the 60th and 80th epochs. In the case of $\phi=0.1$, the learning rate is similarly reduced to 10% of its original value at the 150th and 180th epochs. Each method is repeated three times, and the results are presented in Figures 3 and 4. Here the training loss is calculated as $\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i}^{(t)})$. It is important to note that this selection reflects not the convergence properties of the algorithm, but rather whether the algorithm has fallen into the local minimum point for each agent $\mathbf{x}^{\star}_{i}$. In contrast, the testing loss provides a better indication of the convergence performance.
From the simulation results above, we can see that our algorithm outperforms other methods in the classification of CIFAR-10. Specifically, our proposed algorithm exceeds the performance of those that do not utilize momentum acceleration. When $\phi=1$, both our algorithm and DSGT-HB exhibit comparable convergence performance. However, as heterogeneity increases further ($\phi=0.1$), the convergence performance of DSGT-HB deteriorates significantly, even becoming inferior to that of the original algorithm. In contrast, our algorithm continues to demonstrate robust convergence properties. These findings are consistent with our theoretical analysis.